OpenStudy (anonymous):

one pump can drain a pool in 10 min . when second pump is also used, the pool only takes 9 min to drain. How long would it take the second pump to drain if it were the only pump in use. a. 90minutes b. 89minutes c. 4 14/9 minutes d. 1/90 minutes

OpenStudy (amistre64):

how much is "left in" after 1 minute?

OpenStudy (amistre64):

helps if I read the question thru...

OpenStudy (amistre64):

a + b = 9 minutes a = 10 minutes a:10 as a+b:9 9a = 10a + 10b -a = 10b a = -10b ?? does that make sense?

OpenStudy (amistre64):

d... a = 1/10 drained in one minute a+b = 1/9 drained in a minute 1/10 + b = 1/9 per minute b = 1/9 - 1/10 b = 10-9 /90 b = 1/90 per minute